# ZMP-HH/24-01 The Standard Model from String Theory: What Have We Learned?

### 03 Jan 2024 v1

@@ Abstract
Amidst all candidates of physics beyond the Standard Model, string theory provides a unique proposal for incorporating gauge and gravitational interactions. In string theory, a four-dimensional theory that unifies quantum mechanics and gravity is obtained automatically if one posits that the additional dimensions predicted by the theory are small and curled up, a concept known as compactification. The gauge sector of the theory is specified by the topology and geometry of the extra dimensions, and the challenge is to reproduce all the features of the Standard Model of Particle Physics from them. We review the state-of-the-art in reproducing the Standard Model from string compactifications, together with the lessons drawn from this fascinating quest. We describe novel scenarios and mechanisms that string theory provides to address some of the Standard Model puzzles, as well as the most frequent signatures of new physics that could be detected in future experiments. We finally comment on recent developments that connect, in a rather unexpected way, the Standard Model with Quantum Gravity, and that may change our field theory notion of naturalness.
@@ 2 THE STANDARD MODEL FROM STRING THEORY
In most of the constructions supersymmetry is recovered at the compactification scale which, depending on the details of the model, varies within the range GeV. Assuming supersymmetry at the compactification scale ensures stability, and in particular the absence of dangerous tachyons.111 We will not discuss here compactifications of non-supersymmetric string theories and recent attempts in the literature, for instance Abel:2015oxa (16), to overcome the challenges they face such as tachyonic instabilities and a string-scale cosmological constant. This requisite restricts the shape of the extra dimensions and simplifies the construction of vacua, by reducing second order equations of motion to first order ones, both in the gravity and gauge sectors of the theory. The simplest of these constructions are based on manifolds of special holonomy, like Calabi–Yau threefolds (CY). These are not the most general backgrounds that one may consider to solve the higher-dimensional equations of motion, but they form a large enough set to display all the possibilities of SM model building discovered so far, and the lessons drawn from them.222A larger set of backgrounds that solve the higher-dimensional equations of motion is given by -structure manifolds Gra_a_2006 (17, 12). From a 4d EFT viewpoint, this larger class of backgrounds can be thought of as implementing potentials for light fields and supersymmetry-breaking terms. A full treatment of this larger category of compactifications is beyond the scope of the present text.
As of today, the different string models that realize the SM can be divided in two classes: Those at weak and strong string coupling. Weakly coupled constructions are based on the five supersymmetric perturbative string theories formulated in 10d. The six extra dimensions correspond to a manifold that is either a CY or its quotient by a discrete symmetry of the theory, namely an orbifold or orientifold quotient. The families of constructions that are more developed in searching for the SM are illustrated in Figure 1. They are:
@@
Families of constructions at strong coupling are:
@@ 2.1 Gauge Group
A remarkable result of string theory is that, alternatively, gauge interactions can be localized in space-time. The simplest realization of this idea is given by D-branes, which are necessary objects to reproduce the SM in Type II compactifications Dixon:1987yp (21). In the SM realization by D-branes, each gauge factor comes from a group localized in a slice of the extra dimensions, more precisely in a -submanifold of a compact six-dimensional manifold where D-branes are wrapped, or bound states of these objects. As a consequence, each of these factors is associated to a homology class of , which is the topological information that describes the D-brane charge from a 10d perspective.
@@
A Type II orientifold model is specified by the compact manifold , the orientifold projection and the resulting negative tension objects (O-planes), and finally the D-brane content. The O-plane and D-brane sectors define a set of homology classes of , which are the topological data that determine the 4d gauge group and the chiral spectrum. Thanks to this simple description one can perform a general analysis of 4d chiral gauge anomalies. One finds that the cancellation of the total D-brane charge (a condition known as RR tadpole cancellation) and a 4d generalization Sagnotti:1992qw (22) of the Green-Schwarz mechanism guarantees the cancellation of chiral anomalies. As a byproduct of the Green-Schwarz mechanism several symmetries are rendered massive via a Stückelberg mechanism, and are perceived at low energies as global ’s that are only broken at the non-perturbative level. This is how one is able to recover the exact SM gauge group from an initial D-brane group, together with (perturbative) SM global symmetries like Ibanez:2001nd (23).
@@
Most of these features generalize beyond the heterotic and D-brane setups. In general, the 4d gauge group is associated with a set of objects in the extra dimensions, like submanifolds and gauge bundles on them. These objects have a charge and a tension from a 10d viewpoint, and both must cancel to ensure anomaly cancellation and stability of the gauge sector of the theory. This statement applies to the full gauge sector, which in the most realistic models involves a SM plus a hidden gauge sector. F-theory Vafa:1996xn (24) compactifications with 7-branes, as reviewed e.g. in Weigand:2018rez (25, 26), form the currently best understood example of such non-perturbative generalizations. In these constructions, the available gauge algebras include the full list of ADE type algebras as well as their non-simply laced cousins. What makes this class of constructions particularly interesting for model building is that F-theory combines the possibility of engineering gauge algebras based on exceptional groups with , known from heterotic strings, with the idea of localization of gauge degrees of freedom, as in perturbative Type II orientifolds. This has been studied systematically beginning with BeasleyHeckmanVafaI (27, 28, 29, 30). Such constructions are dual to compactifications of M-theory on (singular) special holonomy spaces, where the non-abelian gauge degrees of freedom result from M2-branes wrapping vanishing 2-cycles.
@@
Another interesting lesson is how string theory models realize gauge coupling unification. Up to corrections, the magnitude of a 4d gauge coupling is specified by how strongly the gauge interaction is diluted in the extra dimensions. In the case of a gauge interaction from a D-brane, this depends on the string coupling and on the volume (in units of the string length ) of the submanifold of that it wraps, leading to
up to subleading corrections. For heterotic compactifications the relevant volume is that of the entire internal space and the result is suppressed by rather that since the gauge degrees of freedom descend from the closed string sector. Because the shape of translates into a series of SM-neutral fields below the compactification scale – dubbed compactification moduli – the SM gauge couplings are in general field dependent. This together with the fact that the SM can be built from D-branes of different sizes and dimensions shows that gauge coupling unification is not a universal feature of string constructions. However, approximate gauge coupling unification is realized when all the SM gauge interactions come from a common source. Schematically, such models start from an initial large gauge group , like , Spin or , which is then endowed with a gauge-group breaking mechanism that yields the SM gauge group. The new mechanisms that string theory provides for gauge-group breaking compared to standard BSM setups are based on the the fact that, microscopically, arises from a higher-dimensional gauge field. Therefore a non-trivial profile for such a gauge field in the extra dimensions can break to the SM group, just like for the heterotic gauge groups discussed above. More details on setups with gauge coupling unification will be provided in Section 3.2.
@@ 2.2 Chiral Spectrum
A general lesson from building 4d EFTs from string theory is that chiral fermions are localized in the extra dimensions. This feature is particularly manifest in Type IIA intersecting D6-brane models, where chiral fermions are localized at the points of intersection of two three-cycles of the compact manifold Blumenhagen:2005mu (33). In heterotic compactifications on a smooth CY the localization is less dramatic, but it also occurs because a chiral fermion is represented by a non-trivial wavefunction in the extra dimensions, whose profile peaks at some particular point. In intersecting D7-branes and F-theory models, chiral fermion localization is a combination of these two mechanisms, and turns out to be crucial to obtain a realistic pattern of Yukawa couplings.
The net number of chiral fermions (that is, up to vector-like pairs) within a given gauge group representation is a quantity protected by gauge invariance. In string theory models, this net chirality is always related to some topological index of the compactification. In the case of intersecting D6-branes, it is simply the signed intersection number of two three-cycles. In heterotic compactifications on smooth CY with internal gauge bundles, it corresponds to the index of the corresponding Dirac operator Green:1987mn (7, 3). This simple description of the chiral spectrum has led to a general modular strategy to find string theory models that contain the SM gauge group and chiral spectrum and are absent of chiral exotics Marchesano:2022qbx (5). The strategy works particularly well for chiral theories with only 2-index tensor representations in the gauge sector, as it happens in D-brane models, and it is equally well suited for smooth Calabi–Yau compactifications, D-branes at singularities Maharana:2012tu (34) and Gepner orientifold models Dijkstra:2004ym (35). It can also be adapted to other constructions, like F-theory models Lin:2014qga (36) and heterotic compactifications with line bundles Blumenhagen:2005ga (20). The description of massless vector-like pairs is in general more involved, as it requires the knowledge of finer geometric data, see e.g. Bies:2017fam (37) in the context of F-theory. However, in concrete models several techniques are available to get rid of vector like-pairs. These include giving a vev for SM singlets or switching on discrete Wilson lines, internal gauge fluxes or background fluxes.
@@
Certain compactification schemes like D-branes at singularities have an upper bound of three SM families, which comes from the number of (complex) extra dimensions Maharana:2012tu (34). This bound is, however, not observed in other setups, and at present there is no compelling argument of why three generations should be preferred from a string theory viewpoint. What instead can be thought of as a universal feature of string theory models is that light fields generically transform in small representations of the gauge group, compared to the existing possibilities in field theory. This is clear in D-brane models, where the open string spectrum always yields 2-index representations of the gauge group. In more general M and F-theory compactifications higher representations are possible Klevers:2017aku (49), but non-generic Taylor:2019ots (50). Establishing bounds on the maximal possible charge is currently an open problem Raghuram:2018hjn (51).
@@ 2.3 Yukawa Couplings
with the Kähler potential of the 4d supergravity theory, a cubic term in the superpotential involving three chiral fields, and their Kähler metrics. Just like gauge couplings, each of these quantities depends on compactification moduli: is a holomorphic and a real function of the 4d fields. Even computing the holomorphic piece for compactifications on general Calabi-Yau varieties requires advanced mathematical techniques (see Lerche:2018gvn (52) and references therein). In general, the Kähler metrics for chiral fields is even more difficult to obtain beyond simple setups. A general expression for the Kähler potential including the chiral fields of the compactification remains as an open problem. In fact, it is oftentimes simpler to directly compute the physical Yukawa couplings instead of each the quantities in the rhs of (2.2), a strategy followed for instance in Cremades:2003qj (53, 54).
The basic Yukawa structure that arises from string compactifications is the following. If the coupling results from a point interaction, like for D-branes at singularities or three D-branes intersecting at a point, then the Yukawa is of order one. If the Yukawa is generated by an overlap of three wavefunctions in the extra dimensions, then there can be a suppression due to the separation of the wavefunction peaks from each other and from wavefunction normalization. With these two features and assuming a minimal Higgs sector, one may try to engineer the observed SM flavor hierarchies. Additional ingredients, elaborated on in Section 3.3.2, which generate Yukawa textures involve global symmetries that forbid certain Yukawa couplings at the perturbative level Ibanez:2008my (55), as well as flavor-dependent discrete gauge symmetries Nilles:2012cy (56, 57, 58).
Despite the progress made in computing and understanding Yukawa couplings, it is fair to say that currently there is no general scheme to reproduce flavor hierarchies. It is however important to stress that string models have been constructed that reproduce the observed SM flavor hierarchies. To name one example, a context in which this has been achieved are F-theory GUT models, discussed in more detail in Section 3.2. The starting point is a Yukawa matrix that at tree-level is of rank one, as can be easily engineered because this rank is a topological quantity Cecotti:2009zf (59). The rank is increased to three when non-perturbative effects from a hidden sector are included Marchesano:2009rz (60), naturally leading to one heavy family of quarks and leptons. In the framework of local F-theory GUTs this scenario can be translated into precise computations Cecotti:2010bp (61), which result in realistic fermion masses and mixings for large parameter regions of the MSSM Font:2012wq (62, 63, 64, 65).
@@ 3.1 Local versus Global Models
This has lead to the idea of local string model building, where in a first approach the gauge sector is analyzed by itself, while the coupling to gravitational degrees of freedom or other hidden sectors is achieved at a later step Aldazabal:2000sa (66). Such an approach is particularly promising when there is a clear separation between the particle physics scale and the Planck scale governing the gravitational dynamics.
Specifically, D-branes placed at singularities of Calabi-Yau threefolds Douglas:1996sw (67) give rise to gauge sectors and charged chiral matter in such local models which can get very close to the Standard Model. Promising classes of such singularities are orbifold-type singularities Aldazabal:2000sa (66, 68), toric singularities (e.g. Krippendorf:2010hj (69) and references therein) or more general del Pezzo singularities such as those analyzed in Verlinde:2005jr (70, 71, 72, 73). An account of progress in the direction of local models has been given in Quevedo:2014xia (74). Questions amenable to purely local techniques include the non-abelian part of the gauge algebra, the associated charged chiral spectrum, and the leading contributions to the Yukawa couplings and the gauge couplings.
This set of questions is to be contrasted with the properties of a string vacuum that are sensitive to the global properties of the model, i.e. to the embedding of the localized degrees of freedom into the compactification manifold. In fact, one of the hallmarks of string theory is that it generically describes gauge and gravitational degrees of freedom not in isolation. Hence special care must be taken in any attempt to separate (certain) particle physics aspects from the quest of coupling the theory back to gravity. As a general lesson, local mechanisms oftentimes require substantial modifications once gravitational effects are taken into account. This applies, in particular, to the sector of abelian continuous Buican:2006sn (72) or of discrete gauge symmetries, to which we will return in Section 3.3.2, to the perturbative Conlon:2009kt (75, 76) or non-perturbative corrections to the Yukawa couplings and gauge kinetic terms, and to all questions related to moduli stabilization, supersymmetry breaking, the stabilization of physical scales such as the gauge coupling unification or supersymmetry breaking scales, as well as the cosmological history, including the evolution of the early universe and late time acceleration. The embedding of (semi-)realistic gauge sectors on D-branes probing local singularities into fully-fledged compact models has been advanced in Cicoli:2012vw (77, 78, 79).

### 08 Feb 2024 v2

# IFT-UAM/CSIC-24-01 ZMP-HH/24-01 The Standard Model from String Theory:

@@ Abstract
Amidst all candidates of physics beyond the Standard Model, string theory provides a unique proposal for incorporating gauge and gravitational interactions. In string theory, a four-dimensional theory that unifies quantum mechanics and gravity is obtained automatically if one posits that the additional dimensions predicted by the theory are small and curled up, a concept known as compactification. The gauge sector of the theory is specified by the topology and geometry of the extra dimensions, and the challenge is to reproduce all of the features of the Standard Model of Particle Physics from them. We review the state-of-the-art in reproducing the Standard Model from string compactifications, together with the lessons drawn from this fascinating quest. We describe novel scenarios and mechanisms that string theory provides to address some of the Standard Model puzzles, as well as the most frequent signatures of new physics that could be detected in future experiments. We finally comment on recent developments that connect, in a rather unexpected way, the Standard Model with Quantum Gravity, and that may change our field theory notion of naturalness.
@@ 2 THE STANDARD MODEL FROM STRING THEORY
In most of the constructions supersymmetry is recovered at the compactification scale, defined as , where is a typical length scale on the compact space. Depending on the details of the model, varies within the range GeV. Assuming supersymmetry at the compactification scale ensures stability, and in particular the absence of dangerous tachyons.111 We will not discuss here compactifications of non-supersymmetric string theories and recent attempts in the literature, for instance Abel:2015oxa (16), to overcome the challenges they face such as tachyonic instabilities and a string-scale cosmological constant. For this reason we will focus on attempts to engineer the appearance of the Minimal Supersymmetric Standard Model (MSSM) as a 4-dimensional Effective Field Theory (EFT) at low energies. This requisite restricts the shape of the extra dimensions and simplifies the construction of vacua, by reducing second order equations of motion to first order ones, both in the gravity and gauge sectors of the theory. The simplest of these constructions are based on manifolds of special holonomy, like Calabi–Yau threefolds (CY). These are not the most general backgrounds that one may consider to solve the higher-dimensional equations of motion, but they form a large enough set to display all of the possibilities of SM model building discovered so far, and the lessons drawn from them.222A larger set of backgrounds that solve the higher-dimensional equations of motion is given by -structure manifolds Gra_a_2006 (17, 12). From a 4d EFT viewpoint, this larger class of backgrounds can be thought of as implementing potentials for light fields and supersymmetry-breaking terms. A full treatment of this larger category of compactifications is beyond the scope of the present text.
As of today, the different string models that realize the SM can be divided in two classes: those at weak and strong string coupling. Weakly coupled constructions are based on the five supersymmetric perturbative string theories formulated in 10d. The six extra dimensions correspond to a manifold that is either a compact CY or its quotient by a discrete symmetry of the theory, namely an orbifold or orientifold quotient. The families of constructions that are more developed in searching for the SM are illustrated in Figure 1. They are:
@@
In some cases, the behavior of string theory at strong coupling is understood through string dualities. This gives additional families of constructions that take advantage of the strong coupling regime. The most developed examples are:
@@ 2.1 Gauge Group
A remarkable result of string theory is that, alternatively, gauge interactions can be localized in space-time. The simplest realization of this idea is given by D-branes, lower-dimensional hypersurfaces in the 10-dimensional space. A D-brane extends along spatial directions plus time. Such objects are necessary to reproduce the SM in Type II compactifications Dixon:1987yp (21). In the SM realization by D-branes, each gauge factor comes from a group localized in a slice of the extra dimensions, more precisely in a -submanifold of a compact six-dimensional manifold where D-branes are wrapped, or bound states of these objects. As a consequence, each of these factors is associated to a homology class of , which is the topological information that describes the D-brane charge from a 10d perspective.
@@
A Type II orientifold model is specified by the compact manifold , the orientifold projection and the resulting fixed planes (O-planes), and finally the D-brane content. The O-plane and D-brane sectors define a set of homology classes of , which are the topological data that determine the 4d gauge group and the chiral spectrum. Thanks to this simple description one can perform a general analysis of 4d chiral gauge anomalies. One finds that the cancellation of the total D-brane charge (a condition known as RR tadpole cancellation) and a 4d generalization Sagnotti:1992qw (22) of the Green-Schwarz mechanism guarantees the cancellation of chiral anomalies. As a byproduct of the Green-Schwarz mechanism several symmetries are rendered massive via a Stückelberg mechanism, and are perceived at low energies as global ’s that are only broken at the non-perturbative level. This is how one is able to recover the exact SM gauge group from an initial D-brane group, together with (perturbative) SM global symmetries like Ibanez:2001nd (23).
@@
Most of these features generalize beyond the heterotic and D-brane setups. In general, the 4d gauge group is associated with a set of objects in the extra dimensions, like submanifolds and gauge bundles on them. These objects have a charge and a tension from a 10d viewpoint, and both must cancel to ensure anomaly cancellation and stability of the gauge sector of the theory. This statement applies to the full gauge sector, which in the most realistic models involves a SM plus a hidden gauge sector. F-theory Vafa:1996xn (24) compactifications with 7-branes, as reviewed e.g. in Weigand:2018rez (25, 26), form the currently best understood example of such non-perturbative generalizations. In these constructions, the available gauge algebras include the full list of ADE type Lie algebras as well as their non-simply laced cousins. What makes this class of constructions particularly interesting for model building is that F-theory combines the possibility of engineering gauge algebras based on exceptional groups with , known from heterotic strings, with the idea of localization of gauge degrees of freedom, as in perturbative Type II orientifolds. This has been studied systematically beginning with BeasleyHeckmanVafaI (27, 28, 29, 30). Such constructions are dual to compactifications of M-theory on (singular) special holonomy spaces, where the non-Abelian gauge degrees of freedom result from M2-branes wrapping vanishing 2-cycles.
@@
Another interesting lesson is how string theory models realize gauge coupling unification. To leading order, the magnitude of a 4d gauge coupling is specified by how strongly the gauge interaction is diluted in the extra dimensions. In the case of a gauge interaction from a D-brane, this depends on the string coupling and on the volume (in units of the string length ) of the submanifold of that it wraps, leading to
up to subleading corrections. For heterotic compactifications the relevant volume is that of the entire internal space and the result is suppressed by rather that since the gauge degrees of freedom descend from the closed string sector.
This gives the compactification-scale boundary condition for the usual renormalization group analysis of couplings in the 4d EFT. The size and shape of is described by the vacuum expectation value of massless scalar fields called compactification moduli. The compactification moduli will translate into SM-singlet fields in the 4d EFT. These have a variety of influences on string constructions, and we will discuss more of their implications below. What is important for us here is that the boundary conditions for different gauge couplings can depend in different ways on these moduli. This together with the fact that the SM can be built from D-branes of different sizes and dimensions shows that gauge coupling unification is not a universal feature of string constructions.
Approximate gauge coupling unification is realized when all of the SM gauge interactions come from a common source. Schematically, such models start from an initial large gauge group , like , Spin or , which is then endowed with a gauge-group breaking mechanism that yields the SM gauge group. The new mechanisms that string theory provides for gauge-group breaking compared to standard BSM setups are based on the the fact that, microscopically, arises from a higher-dimensional gauge field. Therefore a non-trivial profile for such a gauge field in the extra dimensions can break to the SM group, just like for the heterotic gauge groups discussed above. More details on setups with gauge coupling unification will be provided in Section 3.2.
@@ 2.2 Chiral Spectrum
A general lesson from building 4d EFTs from string theory is that chiral fermions are localized in the extra dimensions. This feature is particularly manifest in Type IIA intersecting D6-brane models. In these models chiral fermions are localized at the points of intersection of two three-cycles of the compact manifold Blumenhagen:2005mu (33). In heterotic compactifications on a smooth CY the localization is less dramatic, but it also occurs because a chiral fermion is represented by a non-trivial wavefunction in the extra dimensions, whose profile peaks at some particular point. In intersecting D7-branes and F-theory models, chiral fermion localization is a combination of these two mechanisms. As we will discuss in the next section, chiral fermion localization is crucial to obtain a realistic pattern of Yukawa couplings in this class of constructions.
The number of SM fermions of each type in the 4d EFT is given by the net number of chiral fermion states (that is, up to vector-like pairs) within the given gauge group representation. In string theory models, this net chirality is always related to some topological index of the compactification. In the case of intersecting D6-branes, it is simply the signed intersection number of two three-cycles. In heterotic compactifications on smooth CY with internal gauge bundles, it corresponds to the index of the corresponding Dirac operator Green:1987mn (7, 3). This simple description of the chiral spectrum has led to a general modular strategy to find string theory models that contain the SM gauge group and chiral spectrum and are absent of chiral exotics Marchesano:2022qbx (5). The strategy works particularly well for chiral theories with only 2-index tensor representations in the gauge sector, as it happens in D-brane models, and it is equally well suited for smooth Calabi–Yau compactifications, D-branes at singularities Maharana:2012tu (34) and Gepner orientifold models Dijkstra:2004ym (35). It can also be adapted to other constructions, like F-theory models Lin:2014qga (36) and heterotic compactifications with line bundles Blumenhagen:2005ga (20). The description of massless vector-like pairs is in general more involved, as it requires the knowledge of finer geometric data, see e.g. Bies:2017fam (37) in the context of F-theory. However, in concrete models these vector-like fermions must acquire mass. Several techniques are available for this. These include giving a vev for SM singlets or switching on discrete Wilson lines, internal gauge fluxes or background fluxes.
@@
Certain compactification schemes like D-branes at singularities have an upper bound of three SM families, which comes from the number of (complex) extra dimensions Maharana:2012tu (34). This bound is, however, not observed in other setups, and at present there is no compelling argument of why three generations should be preferred from a string theory viewpoint. What instead can be thought of as a universal feature of string theory models is that the light fields generically transform in small representations of the gauge group, compared to the existing possibilities in field theory. This is clear in D-brane models, where the open string spectrum always yields at most 2-index representations of the gauge group. Note that this is just enough to accommodate the representations occurring in the SM. In more general M and F-theory compactifications higher representations are possible Klevers:2017aku (49), but non-generic Taylor:2019ots (50), with the current record held by Li:2022vfj (51). Establishing bounds on the maximal possible charge is an open problem Raghuram:2018hjn (52).
@@ 2.3 Yukawa Couplings
with the Kähler potential of the 4d supergravity theory, a cubic term in the superpotential involving three chiral fields, and their Kähler metrics. Just like gauge couplings, each of these quantities depends on compactification moduli: is a holomorphic and a real function of the 4d fields. Even computing the holomorphic piece for compactifications on general Calabi-Yau varieties requires advanced mathematical techniques (see Lerche:2018gvn (53) and references therein). In general, the Kähler metrics for chiral fields is even more difficult to obtain beyond simple setups. A general expression for the Kähler potential including the chiral fields of the compactification remains as an open problem. In fact, it is oftentimes simpler to directly compute the physical Yukawa couplings instead of each the quantities in the rhs of (2.2), a strategy followed for instance in Cremades:2003qj (54, 55).
The basic Yukawa structure that arises from string compactifications is the following. If the coupling results from a point interaction, like for D-branes at singularities or three D-branes intersecting at a point, then the Yukawa is of order one. If the Yukawa is generated by an overlap of three wavefunctions in the extra dimensions, then there can be a suppression due to the separation of the wavefunction peaks from each other and from wavefunction normalization. With these two features and assuming a minimal Higgs sector, one may try to engineer the observed SM flavor hierarchies. Additional ingredients, elaborated on in Section 3.3.2, which generate Yukawa textures involve global symmetries that forbid certain Yukawa couplings at the perturbative level Ibanez:2008my (56), as well as flavor-dependent discrete gauge symmetries Nilles:2012cy (57, 58, 59).
Despite the progress made in computing and understanding Yukawa couplings, it is fair to say that currently there is no general scheme to reproduce flavor hierarchies. It is however important to stress that string models have been constructed that reproduce the observed SM flavor hierarchies. To name one example, a context in which this has been achieved are F-theory GUT models, discussed in more detail in Section 3.2. The starting point is a Yukawa matrix that at tree-level is of rank one, as can be easily engineered because this rank is a topological quantity Cecotti:2009zf (60). The rank is increased to three when non-perturbative effects from a hidden sector are included Marchesano:2009rz (61), naturally leading to one heavy family of quarks and leptons. In the framework of local F-theory GUTs this scenario can be translated into precise computations Cecotti:2010bp (62), which result in realistic fermion masses and mixings for large parameter regions of the MSSM Font:2012wq (63, 64, 65, 66).
@@ 3.1 Local versus Global Models
This has lead to the idea of local string model building, where in a first approach the gauge sector is analyzed by itself, while the coupling to gravitational degrees of freedom or other hidden sectors is achieved at a later step Aldazabal:2000sa (67).
Specifically, D-branes placed at singularities of Calabi-Yau threefolds Douglas:1996sw (68) give rise to gauge sectors and charged chiral matter in such local models which can get very close to the Standard Model. Promising classes of such singularities are orbifold-type singularities Aldazabal:2000sa (67, 69), toric singularities (e.g. Krippendorf:2010hj (70) and references therein) or more general del Pezzo singularities such as those analyzed in Verlinde:2005jr (71, 72, 73, 74). An account of progress in the direction of local models has been given in Quevedo:2014xia (75). Questions amenable to purely local techniques include the non-Abelian part of the gauge algebra, the associated charged chiral spectrum, and the leading contributions to the Yukawa couplings and the gauge couplings.
This set of questions is to be contrasted with the properties of a string vacuum that are sensitive to the global properties of the model, i.e. to the embedding of the localized degrees of freedom into the compactification manifold. In fact, one of the hallmarks of string theory is that it generically describes gauge and gravitational degrees of freedom not as being in isolation. Hence special care must be taken in any attempt to separate (certain) particle physics aspects from the quest of coupling the theory back to gravity. As a general lesson, local mechanisms oftentimes require substantial modifications once gravitational effects are taken into account. This applies, in particular, to the sector of Abelian continuous Buican:2006sn (73) or of discrete gauge symmetries, to which we will return in Section 3.3.2, to the perturbative Conlon:2009kt (76, 77) or non-perturbative corrections to the Yukawa couplings and gauge kinetic terms, and to all questions related to moduli stabilization, supersymmetry breaking, the stabilization of physical scales such as the gauge coupling unification or supersymmetry breaking scales, as well as the cosmological history, including the evolution of the early universe and late time acceleration. The embedding of (semi-)realistic gauge sectors on D-branes probing local singularities into fully-fledged compact models has been advanced in Cicoli:2012vw (78, 79, 80).